### v2

# Bayesian Speech Synthesizers Can Learn from Multiple Teachers

Abstract
Text-to-Speech (TTS) is inherently a “one-to-many” mapping characterized by intrinsic uncertainty, yet current paradigms often oversimplify it into a deterministic regression task. While continuous-valued autoregressive (AR) models have recently emerged as a promising alternative to discrete codec-based approaches, they typically rely on a fixed-variance prior, fundamentally constraining generation to a static point estimate that ignores the dynamic variability of natural speech. To bridge this gap, we propose BELLE (Bayesian evidential learning with language modelling), a framework that shifts from deterministic prediction to principled Bayesian inference without increasing model parameters or inference latency. By modeling the acoustic target as a Normal-Inverse-Gamma distribution, BELLE captures data-dependent aleatoric uncertainty. To enable accurate variance estimation on standard single-reference datasets, we introduce a “one-to-many” training strategy that leverages synthetic samples as a statistical support set, allowing the model to learn robust distributional properties rather than merely imitating teacher artifacts. Experiments demonstrate that BELLE, trained on only $\sim$5k hours of data, outperforms leading open-source models trained on 50k hours (achieving a 25.8% relative WER reduction) and naturally supports high-quality streaming generation. Audio samples are available at https://belletts.github.io/Belle/.
TTS, EDL
1 Introduction
Text-to-Speech (TTS) synthesis is fundamentally a “one-to-many” mapping problem. In natural human speech, intrinsic uncertainty is ubiquitous: the same textual content can be spoken with substantial variations in prosody, rhythm, and acoustic details, even by the same speaker. Consequently, a robust TTS system should not merely predict a single deterministic output but rather model the conditional probability distribution of potential speech realizations. However, this stochastic nature is frequently overlooked by the research community, a blind spot reflected in both data availability and modeling paradigms. On the data side, standard TTS datasets predominantly adhere to a rigid “one-to-one” mapping—providing only a single acoustic recording per text—which effectively masks the intrinsic distributional variance of speech. Parallel to this, most existing models treat synthesis as a deterministic regression task or rely on heuristic sampling methods (e.g., top-$p$ or dropout) that lack a principled theoretical basis for modeling data-dependent uncertainty.
In parallel, the architecture of TTS models has evolved rapidly to improve generation fidelity. Autoregressive (AR) models based on discrete audio codecs have gained dominance, exemplified by AudioLM (Borsos et al., 2023) and VALL-E (Wang et al., 2023). While successful, these discrete approaches suffer from quantization errors and information loss inherent to tokenization. To address this, continuous-valued prediction has emerged as a promising alternative (Meng et al., 2024; Liu et al., 2024; Lin and He, 2025; Chen et al., 2024; Wang et al., 2025a), eliminating quantization artifacts by directly modeling continuous acoustic features (e.g., Mel-spectrograms). Despite the shift from discrete to continuous representations improves the representation of speech, it does not by itself address the missing probabilistic modeling of “one-to-many” variability, and probabilistic modeling in this domain remains stagnant. State-of-the-art models like MELLE (Meng et al., 2024) typically rely on a fixed-variance prior assumption. While effective for stability, this formulation fundamentally constrains the model to a point estimate, ignoring the complex, data-dependent aleatoric uncertainty inherent in expressive speech.
To bridge this gap, we propose BELLE (Bayesian evidential learning with language modelling) based on Evidential Deep Learning (EDL) (Amini et al., 2020). Unlike superficial architectural modifications, BELLE represents a substantive shift in modeling philosophy: moving from deterministic regression to principled Bayesian inference. Specifically, BELLE models each acoustic target with a Normal-Inverse-Gamma (NIG) distribution, dynamically predicting variance $\sigma^{2}$ conditioned on context. This capability enables the model to effectively capture the intrinsic “one-to-many” uncertainty of speech, facilitating grounded and diverse sampling that traditional regression objectives cannot achieve. Crucially, this Bayesian treatment integrates seamlessly into the architecture—improving robustness and naturalness without adding any parameters or computational overhead to the inference process compared to standard regression baselines.
Implementing this Bayesian framework, however, introduces a statistical challenge: accurately estimating the variance of a distribution theoretically requires multiple observations for a given input, whereas standard TTS datasets typically provide only a single recording per text. To resolve this, we augment the training set with synthetic samples from diverse pre-trained TTS systems. Crucially, these samples serve not as targets for imitation, but as a statistical support set to characterize the variance structure. By aggregating diverse realizations, our evidential framework learns robust distributional properties, effectively mitigating the influence of potential errors or artifacts from individual teacher models.
[…]
2 Related Work
Continuous-valued Zero-Shot TTS and Streaming TTS.
Continuous-valued TTS models are primarily categorized into autoregressive (AR) and non-autoregressive (NAR) approaches. Among AR methods, MELLE (Meng et al., 2024) directly predicts mel spectrogram frames using a Transformer-based model with Gaussian sampling. Several recent studies have also explored continuous modeling by incorporating diffusion or flow-matching modules into AR frameworks (Wang et al., 2025a; Liu et al., 2024; Jia et al., 2025), or by leveraging VAE-based latent distributions (Lin and He, 2025; Zhu et al., 2024). In contrast, NAR models like E2 TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2024) generate mel-spectrograms via iterative mask-based refinement. Regarding streaming TTS, recent systems (Du et al., 2024; Yang et al., 2024; Sun et al., 2025; Sheng et al., 2025; Wang et al., 2025b) typically combine discrete language models with flow-matching decoders. However, such hybrid architectures often introduce higher complexity and potential latency compared to purely autoregressive solutions.
Evidential Learning vs. Bayesian Neural Networks.
Bayesian modeling in deep learning is broadly divided into approaches that place priors over network weights—classic Bayesian Neural Networks (BNNs) (MacKay, 1992; Neal, 1996)—and those that place priors over the data likelihood parameters, known as Evidential Deep Learning (EDL) (Sensoy et al., 2018; Amini et al., 2020). Standard BNNs typically approximate the weight posterior via Variational Inference (Graves, 2011; Blundell et al., 2015) or Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016). While rigorous, these methods inherently require multiple forward passes or ensemble sampling to estimate predictive uncertainty, introducing significant latency that is prohibitive for real-time streaming TTS. In contrast, EDL adopts an Empirical Bayes (or Type-II Maximum Likelihood) framework. Rather than sampling weights, EDL treats the hyperparameters of the likelihood distribution (e.g., the Normal-Inverse-Gamma parameters) as the prediction targets, allowing the model to analytically marginalize over aleatoric uncertainty in a single deterministic forward pass. This distinction is critical for BELLE: it enables the rigorous modeling of the heteroscedastic, “one-to-many” nature of speech (as detailed in Appendix B) without the computational overhead of weight sampling, effectively balancing high-fidelity probabilistic modeling with the efficiency required for low-latency generation.
3 Mel-based Autoregressive TTS
3.1 Problem Formulation
AR TTS models based on mel-spectrogram synthesis typically formulate the generation process as a sequential next-frame prediction task. Formally, given an input text sequence $\bm{x}=[x_{1},x_{2},...,x_{N}]$, it’s aimed at generating an acoustic mel-spectrogram sequence $\bm{y}=[\bm{y}_{1},\bm{y}_{2},...,\bm{y}_{T}]$, where each frame $\bm{y}_{t}\in\mathbb{R}^{D}$ denotes the spectral representation at time step $t$, and $D$ represents the number of mel-frequency bands. In AR modeling, each mel-spectrogram frame $\bm{y}_{t}$ is generated conditionally depending on the textual content $\bm{x}$ and previous frames $\bm{y}_{<t}=[\bm{y}_{1},...,\bm{y}_{t-1}]$. Thus, the generative process can be described by the following conditional probability decomposition:
$p(\bm{y}|\bm{x};\theta)=\prod\nolimits_{t=1}^{T}p(\bm{y}_{t}|\bm{y}_{<t},\bm{x};\theta)$ (1)
where the conditional probability is typically modelled by a deep learning model with parameters $\theta$.
The streaming generation task is defined that both the input text and target mel-spectrogram are segmented into $M$ consecutive chunks $\bm{X}=[\bm{x}^{(1)},\ldots,\bm{x}^{(M)}]$ and $\bm{Y}=[\bm{y}^{(1)},\ldots,\bm{y}^{(M)}]$. At step $m$, the current audio chunk $\bm{y}^{(m)}$ is predicted conditioned on all previously generated audio chunks $\bm{y}^{(<m)}$ and all available text chunks $\bm{x}^{(\leq m)}$:
$p(\bm{Y}\mid\bm{X};\theta)=\prod_{m=1}^{M}p\big(\bm{y}^{(m)}\mid\bm{y}^{(<m)},\bm{x}^{(\leq m)};\theta\big).$ (2)
3.2 General Architecture of Mel-based Autoregressive TTS Model
Figure 1: The structure of BELLE and the detailed sampling module. The output is assumed to follow a Normal-Inverse-Gamma (NIG) distribution, and the Sampling Module predicts four distribution parameters. Sequentially, variance and mean are obtained via Inverse-Gamma sampling and Gaussian sampling, respectively, followed by a final Gaussian sampling step to generate the output $\bm{y}_{t}^{(1)}$.
General mel-based AR TTS model consists of several interconnected modules: a Prenet, an AR LM, a Sampling Module, a Postnet, a Stop-Prediction Module, and a Vocoder, which are shown in Fig. 1. These modules will be introduced in the following sections.
3.2.1 Prenet and Continuous-Valued AR LM
The continuous-valued AR generation model closely resembles standard LM architectures, specifically decoder-only Transformer models that are frequently employed in contemporary large language models (LLMs), which is shown in the left part of Fig. 1. To delineate the input text clearly, a special $\langle\text{BOS}\rangle$ token is added at the beginning of textual sequences, and an $\langle\text{EOS}\rangle$ token is appended at the end. The Prenet then maps discrete textual tokens $\bm{x}$ into continuous embeddings as well as maps mel-spectrogram frames $\bm{y}$ to the hidden dimension of the AR LM. Next, the AR LM accepts a concatenation of textual embeddings and mel-spectrogram embeddings as inputs, producing corresponding hidden representations $\bm{e}_{t}$.
3.2.2 Sampling Module
The Sampling Module introduces critical stochasticity into the generative process. Specifically, the hidden states $\bm{e}_{t}$ from the AR LM are first projected down to the mel-spectrogram dimension. The projected representation is assumed to follow a specified parametric distribution (e.g., a Gaussian distribution). Subsequently, the Sampling Module samples from this distribution, obtaining an intermediate sampled representation denoted as $\bm{z}_{t}$. Afterwards, a multilayer perception (MLP)-based denoising network refines and further processes the sampled representation, and the resulting denoised output is denoted as $\bm{y}_{t}^{(1)}$ as shown in the right part of Fig. 1. The detailed description of our proposed evidential Bayesian sampling method is provided in Sec. 4.
3.2.3 Stop-Prediction Module and Postnet
Discrete-token AR LM typically incorporates an explicit $\langle\text{EOS}\rangle$ token prediction to naturally terminate sequence generation. For continuous-valued AR generation, a dedicated Stop-Prediction Module is required. Following TransformerTTS or SpeechT5 (Li et al., 2019; Ao et al., 2022), this module predicts a scalar value through a single linear layer followed by a sigmoid activation function, resulting in a score between 0 and 1. Generation stops when this normalized score exceeds a predefined threshold.
The Postnet consists of multiple convolutional layers to refine generated acoustic features like methods in (Shen et al., 2018; Li et al., 2019). Given the output of the Sampling Module $\bm{y}_{t}^{(1)}$, the Postnet predicts a residual term, which is subsequently added back to $\bm{y}_{t}^{(1)}$ and produces the refined output $\bm{y}_{t}^{(2)}$. During training, the model employs the teacher-forcing paradigm; at inference, all intermediate $\bm{y}_{t}^{(1)}$ frames are first collected and then passed through the Postnet to obtain the final output $\bm{y}_{t}^{(2)}$.
3.2.4 Streaming Chunk-based Generation
Figure 2: BELLE-stream adopts the same architecture as BELLE. The text and Mel-spectrogram are split into chunks and interleaved to form the input sequence. Typically, the final chunk generates all remaining audio, making it longer than the preceding ones.
As shown in Fig. 2, a streaming generation strategy by segmenting both the input text and target acoustic sequence into fixed-size chunks is adopted. Let the predefined text chunk size be denoted as $S_{\mathrm{text}}$ and the audio chunk size as $S_{\mathrm{audio}}$. The text sequence is first partitioned into chunks $\bm{x}^{(1)},\ldots,\bm{x}^{(M)}$, where the last chunk may contain fewer than $S_{\mathrm{text}}$ tokens. Similarly, the mel-spectrogram is partitioned into $\bm{y}^{(1)},\ldots,\bm{y}^{(M)}$, each containing up to $S_{\mathrm{audio}}$ frames, with the last chunk possibly exceeding $S_{\mathrm{audio}}$ frames.
The generation process interleaves text and audio chunks in the input to the AR LM, maintaining a causal mask over the sequence. By setting a relatively smaller $S_{\mathrm{audio}}$, each audio chunk is generated with sufficient preceding textual context to ensure coherent synthesis. At inference time, chunk-by-chunk audio generation is performed until the last audio chunk, where the Stop-Prediction Module is invoked to decide the termination point of synthesis.
4 Bayesian Evidential Learning
4.1 Preliminary: EDL in Mel-Spectrogram Prediction
EDL is a Bayesian approach for uncertainty quantification in regression tasks by explicitly modelling the posterior distribution of predictions. In our TTS setting, the observed data $\bm{y}_{t}$ denotes the mel-spectrogram frame at time step $t$, where $\bm{y}_{t}\in\mathbb{R}^{D}$.
Each frame is assumed to follow a Gaussian distribution with unknown mean $\bm{\mu}_{t}$ and variance $\bm{\sigma}_{t}^{2}$, with an Normal-Inverse-Gamma (NIG) conjugate prior:
$\begin{gathered}\bm{y}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right)\\ \bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),\end{gathered}$ (3)
where all hyperparameters $\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t}\in\mathbb{R}^{D}$. The constraints are applied element-wise: $\nu_{t,d}>0,\alpha_{t,d}>1,\beta_{t,d}>0,\forall d\in\{1,\ldots,D\}.$
The posterior predictive distribution is then a multivariate Student-$t$ with diagonal scale:
$p(\bm{y}_{t}|\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t})=\mathrm{St}\!\left(\bm{y}_{t};\,\bm{\gamma}_{t},\,\frac{\bm{\beta}_{t}\odot(1+\bm{\nu}_{t})}{\bm{\nu}_{t}\odot\bm{\alpha}_{t}},\,2\bm{\alpha}_{t}\right),$ (4)
where $\odot$ denotes element-wise multiplication.
In practice, the hyperparameters $(\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t})$ are predicted by the sampling module and optimized via the evidential loss:
$\mathcal{L}_{\text{edl}}(\bm{y}_{t})=\mathcal{L}_{\text{NLL}}(\bm{y}_{t})+\lambda\,\mathcal{L}_{\text{R}}(\bm{y}_{t}),$ (5)
where the NLL enforces distributional fit and the regularization term penalizes incorrect evidence. Details about EDL are in Appendix E.
4.2 From Gaussian Sampling to Bayesian Evidential Sampling
In the MELLE framework (Meng et al., 2024), the latent embedding generation is modeled as a stochastic process. At each timestep $t$, the model predicts the parameters of a multivariate Gaussian distribution from the autoregressive hidden representation $\bm{e}_{t}$. Specifically, a linear projection is used to regress the mean $\bm{\mu}_{t}$ and the diagonal log-variance $\log\bm{\sigma}_{t}^{2}$:
$\left[\bm{\mu}_{t},\log\bm{\sigma}_{t}^{2}\right]=\bm{W}_{\text{G}}\bm{e}_{t}+\bm{b}_{\text{G}}.$ (6)
To enable differentiability during training, the latent embedding $\bm{z}_{t}$ is synthesized using the reparameterization trick, expressed as $\bm{z}_{t}=\bm{\mu}_{t}+\bm{\sigma}_{t}\odot\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$.
A critical component of MELLE is its training objective, which employs a Kullback-Leibler (KL) divergence loss to align the predicted distribution $p_{\bm{\theta}}(\bm{z}_{t}|\bm{e}_{t})$ with a target prior distribution $p(\bm{z}_{t})$. MELLE defines the prior distribution centered at the ground-truth acoustic feature $\bm{y}_{t}$ with a fixed unit variance:
$p(\bm{z}_{t})=\mathcal{N}(\bm{y}_{t},\bm{I}).$ (7)
Consequently, the sampling loss is formulated as:
$\mathcal{L}_{\text{KL}}$ $=\sum_{t=0}^{T-1}D_{\text{KL}}\left(\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\bm{I})\parallel\mathcal{N}(\bm{y}_{t},\bm{I})\right)$
$=\frac{1}{2}\sum_{t=0}^{T-1}\left(\|\bm{\mu}_{t}-\bm{y}_{t}\|^{2}_{2}+\|\bm{\sigma}_{t}\|^{2}_{2}-\sum_{i=1}^{D}\log\bm{\sigma}_{t,i}^{2}-D\right),$ (8)
where $D$ denotes the feature dimensionality. This formulation effectively serves a dual purpose: the term $\|\bm{\mu}_{t}-\bm{y}_{t}\|^{2}_{2}$ acts as a regression loss forcing the predicted mean to track the ground truth, while the variance terms regularize the model uncertainty towards unity, preventing the variance from collapsing to zero.
We propose to replace the Gaussian sampling mechanism with a Bayesian evidential approach. Specifically, NIG distribution parameters are predicted from the AR hidden representation $\bm{e}_{t}$ via a linear layer followed by suitable activation functions to constrain each parameter within appropriate numerical ranges:
$\left[\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t}\right]=\text{Activation}(\bm{W}_{\text{NIG}}\bm{e}_{t}+\bm{b}_{\text{NIG}}),$ (9)
where $\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t}\in\mathbb{R}^{D}$, and $\text{Activation}(\cdot)$ is the selected activation function. Here, suitable activation functions (e.g., softplus for $\bm{\nu}_{t},\bm{\beta}_{t}$, and softplus-shifted for $\bm{\alpha}_{t}$) are applied to ensure that they satisfy the numerical constraints of the NIG distribution (See Sec. 4.1).
During sampling, variance $\bm{\sigma}_{t}^{2}$ is first drawn dimension-wise from an Inverse-Gamma distribution parameterized by $\bm{\alpha}_{t}$ and $\bm{\beta}_{t}$. Conditioned on the sampled variance, the mean $\bm{\mu}_{t}$ is subsequently drawn from a Gaussian distribution with mean $\bm{\gamma}_{t}$ and variance ${\bm{\sigma}_{t}^{2}}/{\bm{\nu}_{t}}$. Finally, conditioned on the sampled parameters $\bm{\mu}_{t}$ and $\bm{\sigma}_{t}^{2}$, $\bm{z}_{t}$ is sampled from another Gaussian distribution with mean $\bm{\mu}_{t}$ and variance $\bm{\sigma}_{t}^{2}$. The overall hierarchical sampling procedure is summarized as:
$\begin{gathered}\bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right),\\ \bm{z}_{t}\sim\mathcal{N}\left(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\right).\end{gathered}$ (10)
Theoretical Comparison
The transition to Bayesian evidential learning offers fundamental theoretical advantages over the standard Gaussian objective. While MELLE minimizes KL divergence against a fixed-variance prior ($\mathcal{N}(\bm{y}_{t},\bm{I})$), implicitly assuming homoscedasticity (uniform uncertainty), our framework adopts an Empirical Bayes approach to perform Type-II Maximum Likelihood Estimation by placing a conjugate NIG prior over the likelihood parameters. This allows the model to dynamically estimate variance from evidence, effectively capturing the inherent heteroscedasticity of speech without increasing backbone parameters. Furthermore, marginalizing over this prior yields a Student-$t$ predictive distribution which strictly subsumes the Gaussian as a limiting case, thereby guaranteeing superior expressive capacity while maintaining analytically tractable, closed-form sampling (refer to Appendix B for details).
4.3 Training Objectives
BELLE is trained end-to-end using a composite loss:
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\lambda_{\text{samp}}\mathcal{L}_{\text{samp}}+\lambda_{\text{flux}}\mathcal{L}_{\text{flux}}+\mathcal{L}_{\text{stop}},$ (11)
where $\lambda_{\text{samp}}$ and $\lambda_{\text{flux}}$ balance the contribution of sampling and flux losses.
The regression loss $\mathcal{L}_{\text{reg}}$ ensures predicted mel-spectrograms closely match the ground truth $\bm{y}^{\text{gt}}$. It includes L1 and L2 terms for both coarse predictions $\bm{y}^{(1)}$ and refined predictions $\bm{y}^{(2)}$:
[…]
The sampling loss $\mathcal{L}_{\text{samp}}$ is used as the EDL loss:
$\mathcal{L}_{\text{samp}}(\bm{y}^{\text{gt}})=\mathcal{L}_{\text{edl}}(\bm{y}^{\text{gt}})=\mathcal{L}_{\text{NLL}}(\bm{y}^{\text{gt}})+\lambda\mathcal{L}_{\text{R}}(\bm{y}^{\text{gt}}),$ (13)
whereas MELLE employs a KL loss for sampling.
The spectrogram flux loss $\mathcal{L}_{\text{flux}}$ encourages temporal dynamics, promoting variability between the current predicted distribution location $\bm{\gamma}_{t}$ and previous-frame ground truth $\bm{y}_{t-1}^{\text{gt}}$:
[…]
In contrast, MELLE uses its Gaussian mean $\bm{\mu}_{t}$.
Lastly, the stop prediction loss $\mathcal{L}_{\text{stop}}$ is a binary cross-entropy loss applied to stop logits from the hidden state $\bm{e}_{t}$. Due to class imbalance, the stop frame receives a higher weight of 500.
4.4 “One-to-Many” Training Strategy
A couple of open-sourced TTS models are used to generate multiple audio samples given the texts provided by the dataset. Given a textual input $\bm{x}$, audio samples are synthesized using a set of $N$ external TTS teacher models, resulting in $N$ synthesized mel-spectrograms. Together with the original human-recorded mel-spectrogram from the dataset, a total of $N+1$ corresponding mel-spectrograms are gotten. Denote these ground-truth mel-spectrograms as $\bm{y}^{\text{gt}}_{i},~i=1,2,\dots,N+1$. Each mel-spectrogram is treated as a ground-truth example and calculated an individual loss corresponding to each one. To balance the contributions from multiple teachers and the original dataset recording, the predefined weights $w_{i}$ are assigned to each ground-truth mel-spectrogram.
Let $\mathcal{L}(\bm{y}^{\text{gt}}_{i})$ denote the complete loss computation defined in Eqn. (11) using $\bm{y}^{\text{gt}}_{i}$ as the ground-truth mel-spectrogram. The final overall training loss is computed as a weighted sum:
$\begin{gathered}\mathcal{L}_{\text{one-to-many}}=\sum\nolimits_{i=1}^{N+1}w_{i}\mathcal{L}(\bm{y}^{\text{gt}}_{i}),\\ \quad\text{with}\quad\sum\nolimits_{i=1}^{N+1}w_{i}=1,\quad w_{i}\geq 0.\end{gathered}$ (15)
5 Experimental Setup
5.1 Training Data and Details
The training dataset is derived from the Librispeech (Panayotov et al., 2015) training set and contains audio samples whose duration is from 0.5 second to 14 seconds, creating a final training dataset containing approximately 706 hours of speech. To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, our training data are augmented by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024), IndexTTS (Deng et al., 2025), SparkTTS (Wang et al., 2025c), F5TTS (Chen et al., 2024), MaskGCT (Wang et al., 2024), and XTTS-v2 (Casanova et al., 2024). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. Details of the data processing methodology can be found in Appendix D.1.
BELLE, BELLE-stream and MELLE are trained on the training dataset, following the training strategy of combining various data sources within each batch (see Sec. 4.4). For BELLE-stream, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. Details about model configuration and training could be found in Appendix D.2 and D.3 respectively.
5.2 Evaluation Settings
The zero-shot TTS capabilities of our model is evaluated using the LibriSpeech test-clean subset by an open-source evaluation protocol (Lee, 2024). Specifically, two inference conditions are considered: Continuation, in which the first 3 seconds of an utterance and its corresponding transcription serve as the prompt, and the model synthesizes the continuation of speech thereafter; and Cross-sentence, where a reference utterance and its corresponding transcription from a given speaker is used as a prompt, and then the model generates speech for a different sentence.
For objective evaluation, word error rate (WER) is reported to measure intelligibility and robustness. WER-C and WER-H are evaluated using Conformer(Gulati et al., 2020) and HuBERT(Hsu et al., 2021) based ASR models respectively. Speaker similarity is measured via cosine similarity of extracted speaker embeddings, with SIM-o referencing the original speech prompt and SIM-r referencing the vocoder-reconstructed prompt.
For subjective evaluation, MOS and SMOS scores are obtained via a crowd-sourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. MOS and SMOS is evaluated following the detailed procedure described in Appendix  H.
For diversity evaluation, building on the layer-wise analysis of WavLM(Chiu et al., 2025), it’s found that the middle-layer features contain rich paralinguistic information, which can be leveraged to assess the acoustic characteristics of speech. Frame-level hidden states from layer-13 are extracted for each sample, mean-pooled and L2-normalized, followed by computation of within-group pairwise cosine similarity and L1/L2 distances.
All evaluation details could be found in Appendix D.5.
6 Results and Discussions
6.1 Main Results
Table 1: Comparison of MOS and SMOS for different systems. Results are reported as mean $\pm$ standard deviation calculated across all ratings for each system.
System Ground Truth MaskGCT F5-TTS MELLE BELLE
MOS $4.20\pm 0.89$ $4.12\pm 0.85$ $\mathbf{4.25}\pm 0.81$ $4.02\pm 0.96$ $4.21\pm 0.75$
SMOS $3.81\pm 0.86$ $3.89\pm 0.90$ $3.95\pm 0.91$ $3.80\pm 0.90$ $\mathbf{4.13}\pm 0.78$
Table 2: Comparison of WER (%) and speaker similarity metrics. GT indicates training with only the original LibriSpeech dataset to validate BELLE’s effectiveness without synthetic data; GT + 6 TTS further includes synthesized data from all six TTS models.
System Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
Ground Truth 1.78 2.15 - 0.668 1.78 2.15 - 0.672
RALL-E (Xin et al., 2024) - - - - 2.5 2.8 - 0.49
ELLA-V (Song et al., 2024) * 2.10 2.91 0.340 0.303 7.15 8.90 0.331 0.307
CLAM-TTS(Kim et al., 2024) - 2.36 0.513 0.477 - 5.11 0.538 0.495
VALL-E (Wang et al., 2023) - 3.8 0.508 - - 5.9 0.580 -
MELLE GT 2.15 2.78 0.491 0.452 7.78 8.76 0.610 0.561
BELLE GT 1.79 2.39 0.492 0.457 3.81 5.01 0.620 0.576
MELLE GT + 6 TTS 2.04 2.59 0.526 0.488 3.30 3.83 0.652 0.606
BELLE GT + 6 TTS 1.63 2.13 0.549 0.519 2.45 2.99 0.679 0.641
Subjective Evaluation (MOS & SMOS).
From Table 1, BELLE, F5-TTS, and Ground Truth all achieve MOS scores around 4.2, suggesting that the generated audio is comparable to natural human speech. Notably, BELLE achieves this competitive performance using less than 5,000 hours of data, compared to $\sim$50,000 hours required by baselines like F5-TTS. While F5-TTS attains a marginally higher mean score (4.25)—likely due to its tendency to produce cleaner audio by removing background noise—BELLE demonstrates superior stability, achieving the lowest standard deviation ($\pm$ 0.75) among all systems, including Ground Truth ($\pm$ 0.89). This indicates that BELLE faithfully replicates the acoustic characteristics of the reference audio without the quality fluctuations often observed in other models. Regarding SMOS, despite the relatively lower Ground Truth score caused by inherent timbre variations in the LibriSpeech dataset, BELLE achieves the highest mean speaker similarity (4.13) with the smallest standard deviation ($\pm$ 0.78), significantly outperforming baselines like F5-TTS ($\pm$ 0.91). This confirms that BELLE captures target timbres more reliably and robustly.
Objective Evaluation.
In terms of objective metrics (Table 2), BELLE achieves state-of-the-art performance, surpassing all baseline systems in both Continuation and Cross-Sentence settings. Specifically, BELLE attains the lowest word error rates (WER-C 1.63% and 2.45%) and the highest speaker similarity scores, significantly outperforming strong baselines such as RALL-E and CLAM-TTS. Crucially, under identical training conditions, BELLE consistently outperforms MELLE across all metrics. For instance, in the Cross-Sentence setting, BELLE reduces WER-H by nearly 22% (from 3.83% to 2.99%) and improves SIM-r from 0.652 to 0.679. These results provide compelling empirical evidence that the proposed Bayesian evidential sampling framework delivers superior robustness and fidelity compared to standard Gaussian sampling.
The Role of Synthetic Data
To understand the source of BELLE’s efficiency and robustness, we analyze the distinct contributions of the Bayesian framework and the synthetic data strategy. First, this effectiveness is intrinsic to the framework, not reliant on synthetic data. As shown in Table 2 (GT setting), BELLE trained solely on human speech drastically outperforms MELLE (e.g., 3.81% vs. 7.78% WER-C). This confirms that the EDL formulation effectively captures uncertainty even from standard corpora by exploiting phonemic recurrence. Second, the synthetic data serves a statistical purpose, not an imitative one. Since standard datasets lack the “one-to-many” variations required for variance estimation, we use synthetic samples to characterize the distributional envelope. By aggregating diverse outputs from multiple teachers, our framework learns robust variance properties without overfitting to the artifacts or errors of any single teacher, keeping the core generation quality anchored in the original human recordings.
6.2 Diversity Analysis
Table 3: Comparison of diversity metrics (cosine distance, L1, and L2) under the Cross-sentence setting. Each of 1,234 LibriSpeech test-clean prompts is generated three times; metrics are computed as within-prompt pairwise distances using WavLM-Large (layer 13) embeddings. BELLE ($\beta*2$) doubles the sampling $\beta$ relative to the default to encourage diversity.
System Cosine distance L1 L2
BELLE ($\beta*2$) 0.0053 3.04 0.095
BELLE 0.0037 2.62 0.080
MELLE 0.0033 2.55 0.078
F5-TTS 0.0005 0.40 0.012
The diversity evaluation is conducted using the evaluation method showed in Sec. 5.2. In the NIG prior, the parameter $\beta_{t}$ denotes the scale of the inverse-gamma distribution over the variance $\sigma_{t}^{2}$. Increasing $\beta_{t}$ raises the expected variance, thereby promoting greater diversity in the sampled outcomes (See Eqn. (10)). From Table 3, it’s observed that BELLE ($\beta*2$) exhibits the largest distances across all three metrics, confirming that increasing the sampling parameter $\beta$ effectively enhances generation diversity. Default BELLE achieves moderate diversity, slightly higher than MELLE, suggesting that Bayesian sampling yields superior diversity compared to Gaussian sampling. In contrast, F5-TTS produces consistently small distances, indicating that its outputs are more deterministic and less diverse across repeated sampling. These findings highlight that, beyond intelligibility and speaker similarity, BELLE offers a controllable trade-off between stability and diversity at inference time, with the $\beta$ parameter serving as a practical knob to adjust sample variability, analogous to the role of the temperature parameter in the sampling process of token-based language models.
6.3 The Effect of Teacher Counts
Table 4: Comparison of WER (%) and speaker similarity metrics for BELLE under different training configurations on a smaller-scale dataset. GT indicates training with only the original LibriSpeech dataset; GT + 2 TTS additionally includes synthesized data from XTTS-v2 and MaskGCT; GT + 6 TTS further includes synthesized data from all six TTS models. data-aug randomly samples from the combined pool of all data sources (GT and 6 TTS) without the evidential “one-to-many” alignment.
System Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
BELLE GT 3.10 3.92 0.344 0.303 6.11 7.34 0.374 0.330
BELLE GT + 2 TTS 2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
BELLE GT + 6 TTS 1.96 2.57 0.444 0.409 4.17 4.73 0.485 0.449
BELLE data-aug 2.10 2.64 0.439 0.404 4.83 5.39 0.480 0.443
Due to limited computational resources, the analysis experiments are conducted on a smaller-scale dataset. The GT setting trains solely on the original LibriSpeech data (Ground Truth) without synthesized augmentation. The GT + 2 TTS setting incorporates additional synthesized speech from two TTS models, while the GT + 6 TTS setting augments the training data with speech from all six TTS models described earlier. The data-aug configuration serves as a baseline using conventional data augmentation, randomly sampling from all available data sources during training. Details about the small dataset and model training configuration can be found in Appendix D.4.
As shown in Table 4, the performance of BELLE consistently improves as the diversity of data sources increases. Furthermore, we exclude the possibility that the performance gain is solely due to the increased volume of training data. In the data-aug setting, the total training data volume is strictly larger than that of the GT + 2 TTS configuration; however, its performance is inferior to GT + 2 TTS on multiple metrics. This degradation may be attributed to the conventional data augmentation strategy, in which the same text can correspond to acoustically different speeches in a random manner, potentially causing confusion for the model. In contrast, under our proposed “one-to-many” training scheme, expanding from GT to GT + 6 TTS leads to consistent performance gains, thereby demonstrating the effectiveness of the evidential training strategy.
Notably, our synthetic multi-sample speech is introduced to provide multiple realizations of the same text—required by the EDL-based Bayesian formulation for uncertainty estimation—rather than to replace human recordings or to imitate any specific teacher model. The core speech generation capability is still grounded in the real human speech (GT), while the synthetic samples act only as statistical observations to characterize conditional variance.
6.4 Results of BELLE-stream
Table 5: Comparison of WER (%) and speaker similarity metrics in the streaming setting.
System Cross-Sentence
WER-C WER-H SIM-r SIM-o
SMLLE(Sun et al., 2025) 5.14 6.37 0.516 0.489
IST-LM(Yang et al., 2024) - 4.53 - 0.653
BELLE-stream 3.54 4.44 0.524 0.491
We evaluate BELLE in a streaming setup to demonstrate its practical superiority in modern conversational pipelines. While we do not propose a novel streaming architecture, this evaluation serves a critical purpose: it validates that our Bayesian evidential framework is orthogonal to and fully compatible with standard autoregressive streaming. Unlike diffusion-based TTS models, which typically suffer from high algorithmic latency due to iterative denoising, BELLE retains the ability to synthesize “on-the-fly” without compromising the generation quality. This confirms that our approach effectively solves the trade-off between high-quality probabilistic modeling and the low Time-to-First-Audio (TTFA) required for duplex dialogue systems.
Table 5 presents the quantitative comparison. BELLE-stream achieves the lowest WER among the streaming systems (3.54%), significantly outperforming SMLLE and IST-LM. Crucially, there is no appreciable degradation in WER compared to the non-streaming BELLE, confirming that evidential modeling imposes no penalty on streaming performance.
In terms of efficiency, BELLE-stream achieves a real-time factor (RTF) of 0.55 for a 10-second utterance. With a chunk size of 0.8 seconds, this results in a first-packet latency (FPL) of approximately 440ms, demonstrating its suitability for real-time applications.
7 Conclusion
In this work, we introduced BELLE, a continuous autoregressive TTS framework that marks a fundamental shift from deterministic point estimation to principled Bayesian inference. By replacing fixed-variance priors with a dynamic Normal-Inverse-Gamma distribution, BELLE explicitly models the intrinsic “one-to-many” aleatoric uncertainty of human speech. To enable robust variance estimation on standard single-reference datasets, we proposed a novel training strategy that leverages synthetic samples not as imitation targets, but as a statistical support set to characterize the solution space.
Empirical results demonstrate that this evidential approach yields significant gains in robustness and naturalness: BELLE trained on $\sim$5k hours of data outperforms leading open-source models trained on 50k hours, achieving a 25.8% relative reduction in WER and superior speaker similarity. Crucially, these improvements are achieved without adding model parameters or inference latency, and the framework naturally extends to high-quality streaming scenarios.
Impact Statement
BELLE introduces a novel Bayesian evidential sampling approach within continuous-valued autoregressive text-to-speech, significantly enhancing the zero-shot TTS synthesis quality. By effectively modeling and generating natural, expressive, and intelligible speech with limited reference data, BELLE advances the flexibility and realism of synthesized audio, making it valuable for various beneficial applications in society. Notably, BELLE can facilitate natural human-machine conversational systems, assistive communication technologies for individuals with speech impairments, and personalized education platforms, thereby positively impacting accessibility, education, and user experiences in interactive dialogue systems.
However, alongside the benefits, zero-shot text-to-speech technologies like BELLE also pose potential risks. They could be misused in unethical or harmful scenarios, such as impersonation, identity fraud, and targeted social engineering attacks. In principle, BELLE could mimic any person’s voice from minimal audio samples, leading to malicious applications aimed at deceiving or misleading individuals. Therefore, responsible use and appropriate safeguards, such as speaker verification, synthetic audio detection, and strong regulatory oversight, are critical directions for addressing and mitigating these societal risks.
Appendix A Limitations
Although BELLE demonstrates strong performance, we acknowledge several limitations in the current study.
First, the current work focuses solely on English speech data without validating multilingual generalization. Extending BELLE to diverse languages and verifying its cross-lingual capabilities remains an important direction for future work.
Second, regarding the experimental setup, we note that the official implementation of MELLE is not publicly available. Consequently, we rely on a faithful re-implementation as the baseline. While potential discrepancies in preprocessing pipelines (e.g., phoneme alignment) may cause the absolute performance of our MELLE baseline to differ slightly from the original paper, the comparison remains rigorously fair and valid, as both BELLE and the MELLE baseline utilize an identical training and evaluation pipeline in our experiments.
Third, due to computational resource constraints, the ablation studies (Table 4 and Table 8) were conducted on a reduced-scale dataset. While the absolute metric values in these studies are not directly comparable to the full-scale results (Table 2), this standard academic practice effectively isolates and verifies the relative contributions of individual components within the proposed framework.
Fourth, we acknowledge that our training data scale (under 5,000 hours) is considerably smaller than that of recent large-scale systems such as F5-TTS and MaskGCT (approx. 50,000 hours). However, this constraint effectively highlights a key strength of BELLE: it achieves comparable or even superior performance to these data-intensive baselines using an order of magnitude less data, demonstrating the exceptional data efficiency of our Bayesian evidential framework.
Finally, as with most autoregressive models, BELLE’s Real-Time Factor (RTF) is higher than that of recent non-autoregressive approaches. Further research on predicting multiple mel-spectrogram frames per step could substantially reduce RTF and enable even lower-latency streaming synthesis.
Appendix B Theoretical Clarification on the Bayesian Framework
In this section, we provide a deeper theoretical discussion regarding the classification of BELLE within the Bayesian Deep Learning landscape. While some strict definitions of Bayesian Neural Networks (BNNs) focus exclusively on placing priors over network weights (weight uncertainty), BELLE follows the Deep Evidential Learning (EDL) paradigm (Amini et al., 2020), which is mathematically grounded in Empirical Bayes (or Type-II Maximum Likelihood). We clarify the rigorous Bayesian nature of our approach and its utility in three aspects.
B.1 Empirical Bayes and Type-II Maximum Likelihood
BELLE does not treat the network weights as random variables, but instead treats the distributional parameters of the data likelihood as random variables. By placing a conjugate Normal-Inverse-Gamma (NIG) prior over the Gaussian likelihood parameters $(\mu,\sigma^{2})$, we formulate the training objective as maximizing the marginal likelihood (or model evidence):
$\log p(\bm{y}\mid\bm{x})=\log\int p(\bm{y}\mid\bm{\theta})p(\bm{\theta}\mid\bm{x})\,d\bm{\theta},$ (16)
where $\bm{\theta}=(\mu,\sigma^{2})$. This process analytically marginalizes over the aleatoric uncertainty parameters, effectively performing Type-II Maximum Likelihood Estimation. This distinguishes BELLE from standard regression (which optimizes point estimates) and firmly positions it within the Bayesian framework of evidence optimization.
B.2 Posterior Sampling in Generative Inference
Contrary to the view that the model “lacks posterior sampling,” BELLE explicitly performs sampling from the predicted posterior during the inference phase. Specifically, the network predicts the hyperparameters $\bm{m}=(\bm{\gamma},\bm{\nu},\bm{\alpha},\bm{\beta})$ of the NIG posterior distribution. The generative process then involves a hierarchical sampling procedure (as defined in Eq. 10):
  1. 1.
    Variance Sampling: We first sample the variance $\bm{\sigma}^{2}$ from the Inverse-Gamma posterior $\Gamma^{-1}(\bm{\alpha},\bm{\beta})$.
  2. 2.
    Mean Sampling: Conditioned on this variance, we sample the mean $\bm{\mu}$ from the Normal distribution $\mathcal{N}(\bm{\gamma},\sigma^{2}/\nu)$.
This stochastic sampling is critical: it breaks the deterministic “one-to-one” mapping typical of standard regression models, allowing BELLE to model the intrinsic “one-to-many” nature of human speech generation.
B.3 The Utility of Uncertainty: Addressing Oversmoothing
Finally, the learned uncertainty serves a crucial role in decision-making within the context of generative modeling. Standard regression approaches (based on MSE) tend to output the “average” of all possible prosodies to minimize error, leading to the well-known “oversmoothing” effect in TTS. In contrast, by explicitly modeling the higher-order uncertainty, BELLE identifies regions of high acoustic ambiguity (e.g., rapid transitions or expressive intonations) and dynamically allocates probability mass. This allows the model to sample valid, distinct acoustic realizations rather than collapsing to the mean, directly resulting in the enhanced prosodic diversity and naturalness observed in our experiments.
B.4 Theoretical Comparison with Standard Gaussian Sampling
Beyond the classification of the framework, we provide a rigorous theoretical comparison demonstrating why the proposed Evidential Deep Learning (EDL) approach is superior to the standard Gaussian sampling used in baseline models like MELLE.
Expressive Capacity (Student-$t$ as a Superset)
Theoretically, the EDL framework offers strictly greater expressive capacity. By marginalizing over the conjugate Normal-Inverse-Gamma (NIG) prior, the posterior predictive distribution follows a Student-$t$ distribution (Amini et al., 2020). Since the Gaussian distribution is mathematically defined as the limiting case of the Student-$t$ distribution as the degrees of freedom (represented by the evidence parameter $\nu$) approach infinity, our model’s capacity inherently subsumes that of the standard Gaussian approach. This ensures that BELLE retains the ability to model standard Gaussian distributions while providing the crucial additional flexibility to capture heavy-tailed distributions and complex outliers in the mel-spectrogram space.
Dynamic Variance vs. Fixed Regularization
A fundamental limitation of MELLE’s Gaussian sampling lies in its training objective. As shown in Eq. (4.2), MELLE minimizes the KL divergence against a prior distribution with a fixed unit variance ($\mathcal{N}(\bm{y}_{t},\bm{I})$). This term acts as a regularizer that forces the predicted variance towards $\bm{I}$, implicitly imposing a homoscedastic assumption (i.e., treating uncertainty as uniform across all speech frames). In contrast, our Bayesian approach enables dynamic variance estimation. By utilizing Type-II Maximum Likelihood, the model learns to estimate aleatoric uncertainty directly from the input evidence without the inductive bias of a fixed prior. This allows BELLE to effectively model the inherent heteroscedasticity of speech—dynamically assigning higher variance to ambiguous pronunciations and lower variance to deterministic segments—thereby improving robustness and generation quality.
Appendix C Modeling Temporal and Spectral Dependencies
A potential theoretical concern regarding the proposed framework is the assumption of a diagonal covariance matrix in the output distribution (i.e., independent Gaussian/NIG distributions per log-mel bin). Critics might argue that this factorization neglects the complex temporal and cross-band (spectral) correlations inherent in physical speech signals. In this section, we clarify how BELLE effectively captures these dependencies through its autoregressive backbone.
While the output distribution parameters at timestep $t$ are factorized dimension-wise, the hidden representations from which they are derived are deeply contextual and correlated. The model does not predict parameters based on an isolated frame; rather, the prediction is conditioned on a high-dimensional hidden state $\bm{h}_{t}$ generated by the Transformer decoder.
Mathematically, let $\bm{\Phi}_{t}=[\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t}]$ denote the predicted NIG distribution parameters at time $t$. The generation process is governed by:
$\bm{\Phi}_{t}=f_{\text{proj}}(\bm{h}_{t}),$ (17)
where $f_{\text{proj}}$ is the linear projection head. Crucially, the hidden state $\bm{h}_{t}$ is computed autoregressively:
$\bm{h}_{t}=\text{TransformerBlock}(\bm{e}_{<t},\bm{h}_{<t}),$ (18)
where $\bm{e}_{<t}$ represents the history of acoustic tokens.
Therefore, although the projection $f_{\text{proj}}$ maps $\bm{h}_{t}$ to factorized parameters, $\bm{h}_{t}$ itself serves as a compact summary of the entire temporal history and spectral context.
  • Temporal Dependency: Since $\bm{h}_{t}$ encapsulates information from $\bm{h}_{1:t-1}$, the predicted parameters $\bm{\Phi}_{t}$ are strictly conditioned on the temporal trajectory, ensuring physical continuity and coherence over time.
  • Spectral Dependency: The hidden state $\bm{h}_{t}$ is a dense vector where all frequency bands effectively interact within the self-attention mechanism layers before the final projection. Thus, cross-band correlations are implicitly modeled within the latent space before being disentangled at the final output layer.
Consequently, the factorization at the output layer is a design choice for computational tractability and closed-form sampling, which does not compromise the model’s ability to capture complex dependencies encoded in the autoregressive hidden states.
Appendix D Detailed Experimental Setup
D.1 Training Data
The training dataset is derived from the Librispeech (Panayotov et al., 2015) training set. We preprocess this dataset using a voice activity detection (SileroTeam, 2024) algorithm to remove prolonged silent intervals. Subsequently, to ensure quality and manageability, we pick out audio samples whose duration is from 0.5 second to 14 seconds, creating a final training dataset containing approximately 706 hours of speech. All audio samples are resampled to 16 kHz and converted into 80-dimensional mel-frequency spectrograms. Additionally, we apply grapheme-to-phoneme (G2P)  [^1] https://github.com/espeak-ng/espeak-ng conversion to preprocess textual transcriptions.
To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, we augment our training data by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024) (LLM + flow matching-based streaming TTS), IndexTTS (Deng et al., 2025) (AR discrete acoustic token TTS), SparkTTS (Wang et al., 2025c) (LLM-based single-stage TTS), F5TTS (Chen et al., 2024) (flow matching-based efficient TTS), MaskGCT (Wang et al., 2024) (masked generative NAR TTS), and XTTS-v2 (Casanova et al., 2024) (cross-lingual expressive TTS). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. For consistency, all synthesized audio samples generated by these models are also resampled to 16 kHz.
D.2 Model Configurations
Input mel-spectrograms first pass through a 3-layer prenet with dropout of 0.5 in both training and inference stages following Tacotron (Wang et al., 2017). The AR LM is a decoder-only Transformer consisting of 12 blocks, each with 16 attention heads, a hidden size of 1024, a feed-forward dimension of 4096, and a dropout rate of 0.1  [^2] Our code is modified from https://github.com/lifeiteng/vall-e.. The sampling module includes a linear projection to derive the sampling parameters and a 3-layer residual MLP for denoising. A post-processing module comprising five convolutional blocks (a kernel size of 5 and 256 channels) is applied for mel-spectrogram refinement. The final waveform is synthesized using the pretrained HiFi-GAN vocoder (Kong et al., 2020) [^3] The pretrained vocoder can be found in https://huggingface.co/mechanicalsea/speecht5-tts.
D.3 Training Details of Main Results
We train BELLE, BELLE-stream and reproduce MELLE on the data in Sec.5.1, following the training strategy of combining various data sources within each batch (see Sec. 4.4), with predefined training weights that the original Librispeech audio is weighted by 0.22 and each synthesized audio source (six TTS models) is weighted by 0.13, which ensures the original Librispeech audio receives approximately twice the weighting of synthesized audio, reflecting relative considerations of synthetic audio quality. Models are trained by AdamW optimiser with a total batch size of about 160K frames distributed across 16 NVIDIA A800 GPUs. BELLE and MELLE’s training proceeds for 450K updates, where the learning rate is first linearly warmed up to a peak value of $5\times 10^{-4}$ over the initial 10% of training steps and thereafter linearly decayed to zero. Regarding the hyperparameters for BELLE, we set $\lambda=0.5$ in Eqn. (13) and $\lambda_{\text{samp}}=0.2$, $\lambda_{\text{flux}}=0.5$ in Eqn. (11). As for MELLE, our hyperparameter settings strictly follow the original MELLE paper, where $\lambda_{\text{samp}}=0.1$, $\lambda_{\text{flux}}=0.5$.
BELLE-stream is initialized from the trained BELLE model, trained with a batch size of 80K frames across 8 NVIDIA A800 GPUs for 150K updates, using $\lambda_{\text{flux}}=0.1$ while keeping all other hyperparameters identical to BELLE. In our implementation, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. Here, $S_{\mathrm{text}}$ is computed from the number of phonemes obtained after G2P conversion, while $S_{\mathrm{audio}}$ is determined from the number of mel-spectrogram frames. Let $L_{\mathrm{text}}$ and $L_{\mathrm{audio}}$ denote the total phoneme count and the total mel-frame count of an utterance, respectively. During training, we filter out audio samples whose ratio $L_{\mathrm{audio}}:L_{\mathrm{text}}$ is less than $2.5$, ensuring that each text chunk retains sufficient corresponding audio frames for effective streaming generation.
D.4 Training Details of Analysis Results
Due to limited computational resources, the analysis experiments are conducted on a smaller-scale dataset. The training dataset is derived from the Librispeech train-clean 100-hour subset, following the same processing procedure in Sec. 5.1, which contains approximately 72 hours of speech. Training solely on the original Librispeech data without synthesized augmentation is denoted as 1-teacher. The 3-teacher condition introduces additional synthesized speech from XTTS-v2 and MaskGCT, about 218h. The 7-teacher condition includes augmented speech from all six TTS models mentioned in Sec. 5.1, about 455h. The data-aug configuration serves as a strong baseline using a conventional data augmentation strategy, where data from all sources totaling 455 hours is randomly sampled during training. The 3-teacher and 7-teacher follow the training strategy in Sec. 4.4, with a total batch size of about 80K frames and 47K training steps.
D.5 Evaluation Settings
We evaluate the zero-shot TTS capabilities of our model using the LibriSpeech test-clean subset following VALL-E (Wang et al., 2023) by an open-source evaluation protocol (Lee, 2024). Specifically, we consider two inference conditions: Continuation, in which the first 3 seconds of an utterance and its corresponding transcription serve as the prompt, and the model synthesizes the continuation of speech thereafter; and Cross-sentence, where we use a reference utterance and its corresponding transcription from a given speaker as a prompt, and then the model generates speech for a different sentence while preserving speaker characteristics.
For objective evaluation, we report word error rate (WER) to measure intelligibility and robustness, using two automatic speech recognition models: a Conformer-Transducer [^4] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge (Gulati et al., 2020) and a fine-tuned HuBERT-Large model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021). We denote results obtained from these systems as WER-C and WER-H, respectively. To quantify speaker similarity, we calculate the cosine similarity between extracted speaker embeddings using a WavLM-TDCNN model [^6] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022). We provide two similarity metrics: SIM-o computes the similarity against the original speech prompt, whereas SIM-r uses the vocoder-reconstructed version of the prompt.
For subjective evaluation, we obtain MOS and SMOS scores via a crowdsourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. From the audio samples generated under the Cross-sentence setting, we select one sample per speaker, resulting in a total of 40 audio samples. All samples from different systems are resampled to 16kHz to ensure a fair comparison. We evaluate MOS and SMOS following the detailed procedure described in Appendix  H. Note that BELLE-stream is fine-tuned to a fixed speaker timbre and does not use an audio prompt; therefore, it is not evaluated for SMOS or SIM metrics, and its results are reported only under the Cross-sentence setting.
For diversity evaluation, building on the layer-wise analysis of WavLM in (Chiu et al., 2025), we adopt the 13th (of 24) layer of WavLM-Large as the speaker representation; 1,234 audio prompts are sampled from the LibriSpeech test-clean subset, and under the Cross-sentence setting each prompt is inferred three times to obtain 1,234 groups of outputs (three per group); frame-level hidden states from layer 13 are extracted for each generation, mean-pooled and L2-normalized, and within-group pairwise cosine similarity, L1, and L2 distances are computed; corpus-level means and standard deviations across all 1,234 groups are reported to quantify the diversity of the generated speech.
To maintain fairness and consistency across all evaluations, we filter the test utterances to only those between 4 and 10 seconds in duration, and report all evaluation metrics using a fixed evaluation set shared among all compared models.
Appendix E Details about Bayesian Evidential Learning
Bayesian Evidential Learning (EDL) provides a principled framework for uncertainty estimation in regression tasks by explicitly modeling the posterior distribution of predictions. Observed data $y$ are assumed to be drawn independently and identically distributed (i.i.d.) from a Gaussian distribution with unknown mean and variance. According to Maximum Likelihood Estimation (MLE), we aim to find parameters $\mu$ and $\sigma^{2}$ that maximize the likelihood of observing the given data $y$, or equivalently, minimize the Negative Log Likelihood (NLL) loss. We model this problem by placing prior distributions on both the mean and variance:
$y_{1},\ldots,y_{N}\sim\mathcal{N}(\mu,\sigma^{2}),\quad\mu\sim\mathcal{N}\left(\gamma,\frac{\sigma^{2}}{\nu}\right),\quad\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$ (19)
where $\Gamma(\cdot)$ denotes the gamma function, $\bm{m}=(\gamma,\nu,\alpha,\beta)$ summarizes the set of hyperparameters, and the constraints are given by $\gamma\in\mathbb{R}$, $\nu>0$, $\alpha>1$, and $\beta>0$. We set $q(\mu,\sigma^{2})=p(\mu,\sigma^{2}\mid y_{1},\ldots,y_{N})$ as our target posterior. For tractability, we adopt a approximation, assuming $q(\mu,\sigma^{2})=q(\mu)q(\sigma^{2})$. Under this assumption, we use the Normal Inverse-Gamma (NIG) distribution as our conjugate prior, which can be written as:
$p(\mu,\sigma^{2}\,|\,\gamma,\nu,\alpha,\beta)=\frac{\beta^{\alpha}\sqrt{\nu}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\left(\frac{1}{\sigma^{2}}\right)^{\alpha+1}\exp\left\{-\frac{2\beta+\nu(\gamma-\mu)^{2}}{2\sigma^{2}}\right\}.$ (20)
Denote $\bm{\theta}=(\mu,\sigma^{2})$, marginalizing out $\mu$ and $\sigma^{2}$:
$p(y_{i}\mid\bm{m})=\frac{p(y_{i}\mid\bm{\theta},\bm{m})p(\bm{\theta}\mid\bm{m})}{p(\bm{\theta}\mid y_{i},\bm{m})}=\int_{\sigma^{2}=0}^{\infty}\int_{\mu=-\infty}^{\infty}p(y_{i}\mid\mu,\sigma^{2})\,p(\mu,\sigma^{2}\mid\bm{m})\,\mathrm{d}\mu\,\mathrm{d}\sigma^{2}$ (21)
Substituting the Gaussian likelihood and the Normal Inverse-Gamma prior into the above equation yields an analytically tractable form, Student-$t$ distribution:
$p(y_{i}|\gamma,\nu,\alpha,\beta)=\text{St}\left(y_{i};\gamma,\frac{\beta(1+\nu)}{\nu\alpha},2\alpha\right),$ (22)
where $\gamma$ is the location, $\frac{\beta(1+\nu)}{\nu\alpha}$ is the scale, and $2\alpha$ is the degrees of freedom.
In practice, a neural network predicts the NIG parameters directly. The training loss, termed evidential loss, comprises a negative log-likelihood term and a regularization term that penalizes incorrect evidence:
$\mathcal{L}_{\text{EDL}}(y_{i})=\mathcal{L}_{\text{NLL}}(y_{i})+\lambda\mathcal{L}_{\text{R}}(y_{i})$ (23)
where
$\mathcal{L}_{\text{NLL}}(y_{i})$ $=\frac{1}{2}\log\left(\frac{\pi}{\nu}\right)-\alpha\log(\Omega)+\left(\alpha+\frac{1}{2}\right)\log\left(\nu(y_{i}-\gamma)^{2}+\Omega\right)+\log\left(\frac{\Gamma(\alpha)}{\Gamma\left(\alpha+\frac{1}{2}\right)}\right)$ (24)
$\mathcal{L}_{\text{R}}(y_{i})$ $=|y_{i}-\gamma|\cdot(2\nu+\alpha)$ (25)
with $\Omega=2\beta(1+\nu)$.
Appendix F MELLE
F.1 Gaussian Sampling in MELLE
In MELLE (Meng et al., 2024), the latent sampling module assumes that the embedding at timestep $t$, denoted as $\bm{z}_{t}$, is drawn from a multivariate Gaussian distribution:
$\bm{z}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\bm{I}),$ (26)
where the mean $\bm{\mu}_{t}\in\mathbb{R}^{D}$ and the log-variance $\log\bm{\sigma}_{t}^{2}\in\mathbb{R}^{D}$, with $D$ representing the number of mel-frequency bands, are computed from the autoregressive hidden representation $\bm{e}_{t}$ via a linear transformation:
$\left[\bm{\mu}_{t},\log\bm{\sigma}_{t}^{2}\right]=\bm{W}_{\text{G}}\bm{e}_{t}+\bm{b}_{\text{G}}.$ (27)
Sampling is performed using the standard reparameterization trick:
$\bm{z}_{t}=\bm{\mu}_{t}+\bm{\sigma}_{t}\odot\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}).$ (28)
F.2 Sampling Loss
When Gaussian sampling is used, the sampling loss is typically implemented as a Kullback–Leibler (KL) divergence loss between the predicted Gaussian distribution and a predefined Gaussian prior distribution with mean $\bm{y}^{\text{gt}}$ and variance $\bm{I}$.
Appendix G Additional Experimental Analysis
In this section, we provide further analysis regarding the hyperparameter choices in the multi-teacher framework and the sampling strategy.
G.1 Weight Allocation in Multi-Teacher Learning
In our multi-teacher learning framework, we assign specific weights to the losses derived from different teacher models. The allocation criteria reflect our prioritization of the original ground-truth audio. Specifically, we aim to ensure that the original LibriSpeech audio receives approximately twice the weight of any individual synthetic audio source. Given 6 synthetic sources, our final configuration assigns a weight of 0.22 to LibriSpeech and 0.13 to each synthetic source (totaling $\approx 1.0$).
To validate this choice, we conducted an ablation study on a reduced-scale dataset (consistent with Sec. 6.3). As shown in Table 6, the configuration with weights 0.22/0.13 yields the optimal balance, particularly minimizing WER. Deviating from this ratio (either increasing or decreasing the weight of real data) leads to a degradation in performance.
Table 6: Ablation study on teacher weight allocation. “Weight (Real)” refers to the original LibriSpeech data, and “Weight (Syn)” refers to each of the 6 synthetic sources.
Weight (Real) Weight (Syn) Continuation Cross-Sentence
WER-H $\downarrow$ SIM-o $\uparrow$ WER-H $\downarrow$ SIM-o $\uparrow$
0.16 0.14 2.64 0.415 5.17 0.453
0.22 0.13 2.57 0.409 4.73 0.449
0.28 0.12 2.78 0.406 5.47 0.444
0.34 0.11 2.76 0.405 5.47 0.446
G.2 Trade-off Between Diversity and Robustness
In Table 3 of the main text, we demonstrated that doubling the sampling parameter $\beta$ (denoted as BELLE ($\beta*2$)) significantly enhances the diversity of the generated speech (higher Cosine distance, L1, and L2 scores). However, in generative modeling, there is typically a trade-off between diversity and robustness/accuracy. To quantify this, we evaluated the WER and SIM metrics for the $\beta*2$ setting.
As shown in Table 7, while scaling $\beta$ improves diversity, it results in a moderate increase in WER (from 2.45% to 3.33%) and a slight decrease in speaker similarity. This indicates that a larger $\beta$ encourages the model to explore the “tails” of the predicted distribution, producing more varied and expressive prosody at the cost of slightly reduced stability. This trade-off is expected and allows users to tune $\beta$ according to their specific application requirements (e.g., favoring stability for reading tasks vs. diversity for creative tasks).
Table 7: Impact of doubling the sampling parameter $\beta$ on performance metrics (Cross-Sentence setting). This should be viewed in conjunction with Table 3 which shows the corresponding increase in diversity.
System WER-C $\downarrow$ SIM-o $\uparrow$
BELLE (Default) 2.45 0.641
BELLE ($\beta*2$) 3.33 0.612
G.3 The effects of the sampling module and flux loss
Table 8: Ablation study examining the effects of the sampling module and flux loss on BELLE GT + 2 TTS. ✓ indicates enabled, ✗ indicates disabled.
Sampling Flux Loss Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
4.30 5.00 0.382 0.344 12.04 13.07 0.392 0.351
2.50 3.25 0.388 0.347 5.61 6.37 0.417 0.372
2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
An ablation study is conducted to investigate how the sampling module and flux loss affect the performance of the BELLE model. Due to limited computational resources, the analysis experiments are also conducted on the smaller-scale dataset. Our experiments are based on the BELLE GT + 2 TTS setting. As demonstrated in Table 8, the sampling module plays a more crucial role than flux loss. Removing the flux loss results in only a slight performance degradation, whereas removing the sampling module leads to a severe performance drop. This indicates that Bayesian sampling plays a crucial role in the effectiveness of BELLE.
Appendix H Subjective Evaluation
(a) Speech Similarity MOS
(b) Speech Quality MOS
Figure 3: Screenshots of subjective evaluations.
For the TTS task, we focus on the Mean Opinion Score (MOS) and Speaker Similarity (SMOS). From the audio samples generated under the Cross-sentence setting, we select one sample per speaker, resulting in a total of 40 audio samples. All samples from different systems are resampled to 16kHz to ensure a fair comparison. The details are as follows: For speech quality evaluation, we conducted an MOS (Mean Opinion Score) test and explicitly instructed the raters to focus on assessing audio quality and naturalness, while ignoring differences in style (e.g., timbre, emotion, and prosody). The raters were presented with and scored samples, and each rater was asked to evaluate the subjective naturalness on a 1-5 Likert scale. When scoring, the order of audio samples is randomly shuffled.
For speaker similarity evaluation, we asked the raters to focus on the similarity of the speaker’s identity (timbre) to the reference, while ignoring differences in content, grammar, or audio quality. We paired each synthetic utterance with a reference utterance to assess the degree of matching between the synthesized speech and the target speaker. Each rater was asked to evaluate the speaker similarity on a 1-5 Likert scale. When scoring, the order of audio samples is randomly shuffled.
Our subjective evaluation was crowd-sourced, with 15 native speakers participating via Amazon Mechanical Turk. The instructions for the testers are shown in the Figure 3. We paid approximately \$100 in participant compensation. A small portion of the speech samples used in the test can be heard at the following website: https://belletts.github.io/Belle/.